# Learning When to Think: Shaping Adaptive Reasoning in R1-Style Models via Multi-Stage RL

### v1

@@ Abstract
Large reasoning models (LRMs) are proficient at generating explicit, step-by-step reasoning sequences before producing final answers. However, such detailed reasoning can introduce substantial computational overhead and latency, particularly for simple problems. To address this over-thinking problem, we explore how to equip LRMs with adaptive thinking capabilities—enabling them to dynamically decide whether or not to engage in explicit reasoning based on problem complexity. Building on R1-style distilled models, we observe that inserting a simple ellipsis ("…") into the prompt can stochastically trigger either a thinking or no-thinking mode, revealing a latent controllability in the reasoning behavior. Leveraging this property, we propose AutoThink, a multi-stage reinforcement learning (RL) framework that progressively optimizes reasoning policies via stage-wise reward shaping. AutoThink learns to invoke explicit reasoning only when necessary, while defaulting to succinct responses for simpler tasks. Experiments on five mainstream mathematical benchmarks demonstrate that AutoThink achieves favorable accuracy–efficiency trade-offs compared to recent prompting and RL-based pruning methods. It can be seamlessly integrated into any R1-style model, including both distilled and further fine-tuned variants. Notably, AutoThink improves relative accuracy by 6.4% while reducing token usage by 52% on DeepSeek-R1-Distill-Qwen-1.5B, establishing a scalable and adaptive reasoning paradigm for LRMs.
Project Page: https://github.com/TU2021/AutoThink.
@@ 1 Introduction
Recently, reasoning-focused Large Language Models (LLMs), also referred to as Large Reasoning Models (LRMs) [37], have demonstrated remarkable progress in solving complex reasoning tasks. Particularly, DeepSeek-R1 [7] uses only outcome-based feedback and incentivizes explicit reasoning capabilities through reinforcement learning (RL) with verifiable rewards. DeepSeek-R1 and its distilled models typically follow the <think> and <answer> format, where the <think> process generates explicit, step-by-step reasoning sequences to support obtaining a final answer during the <answer> phase. We refer to models that follow this Chain of Thought (CoT) [33] prompting scheme as R1-style models. The explicit thinking process, which enables self-reflection, backtracking, and validation, is widely regarded as essential for enhancing reasoning accuracy. Arising from this understanding, a popular paradigm has emerged that improves solution quality by increasing thinking token allocation during inference-time reasoning [15, 36]. However, this paradigm introduces a major bottleneck: excessive thinking token generation leads to high computational cost and latency, raising the over-thinking phenomenon, where many reasoning steps are redundant or inefficient [28, 12].
To mitigate over-thinking, recent efforts have explored hybrid reasoning and concise reasoning strategies. In the industry, Claude 3.7 Sonnet [2] introduces a controllable reasoning framework that allows the model to switch between standard and extended reasoning modes. Similarly, Qwen3 [29] proposes a thinking control scheme with a "thinking" mode (slow thinking) and a "non-thinking" mode (fast thinking), and provides users with the flexibility to choose whether the model should engage in reasoning or not. In the academic community, parallel research has focused on designing prompt-guided efficient reasoning [38, 23] or training pruning-based models to achieve concise reasoning [10, 6, 39]. While promising, these approaches either rely on manually predefined modes or uniformly prune reasoning steps, which may degrade performance on harder instances. A fundamental question then arises to address the over-thinking issue:
@@ 2.1 A Surprising Effect of Minimal Prompt Modification
Recent efforts on concise reasoning aims to eliminate unnecessary thought, either via prompting that explicitly bypasses thinking [18], or RL-based training that penalizes long outputs [6]. While effective at shortening responses, these methods enforce uniform brevity regardless of problem complexity. Rather than compressing by default, we pose a subtler question:
@@ 3.1 Stage 1: Preventing Mode Collapse by Batch Reward Balance
While this reward structure encourages difficulty-aware behavior, it causes instability during early training. The model may collapse into a degenerate policy, either always thinking or always skipping, depending on which yields higher expected reward in the short term. This limits exploration and hinders later optimization. To mitigate this, we introduce batch-level reward balancing:
@@ 3.2 Stage 2: Reinforcing Reliable Behavior within Dual Modes
To allow the model to refine its behavior without external constraints, we remove the batch-level balancing used in previous stage and allow free evolution of the reasoning policy. The reward is set directly to the naive definition:
@@ 4.1 Setup
AutoThink-Stage2 92.2 38.5 56.2 57.1 83.7 65.5 2519 2980 5797 8676 4925 4979 3.0
AutoThink-Stage3 91.2 38.2 56.4 54.8 83.3 64.8 2146 2838 5498 8051 4645 4635 1.1
@@ 4.4 In-Depth Behavioral and Efficiency Analysis
We further conduct a series of additional analyses on reasoning behaviors, training cost, and a case study. Details are presented in Appendix B due to space limitations.
@@ Appendix A Additional Prompt Variants
Appendix B Extended Experimental Results
B.1 Reasoning Behavior Profiling
@@
B.2 Training Cost Comparison
@@
B.3 Case Study
Figure 10,11,12 presents some examples comparing four prompting strategies. For easy problems, AutoThink produces correct answers without explicit reasoning, reflecting effective fast thinking. For medium problems, it may activate both reasoning modes, with thinking and no-thinking responses potentially coexisting. For hard problems, the model engages in deeper, slower reasoning, demonstrating iterative understanding and self-verification before arriving at the correct solution. These observations demonstrate how AutoThink adapts its reasoning to problem difficulty, balancing efficiency and reliability through dynamic control of reasoning depth.

### v2

@@ Abstract
Large reasoning models (LRMs) are proficient at generating explicit, step-by-step reasoning sequences before producing final answers. However, such detailed reasoning can introduce substantial computational overhead and latency, particularly for simple problems. To address this overthinking problem, we explore how to equip LRMs with adaptive thinking capabilities—enabling them to dynamically decide whether or not to engage in explicit reasoning based on problem complexity. Building on R1-style distilled models, we observe that inserting a simple ellipsis ("…") into the prompt can stochastically trigger either a thinking or no-thinking mode, revealing a latent controllability in the reasoning behavior. Leveraging this property, we propose AutoThink, a multi-stage reinforcement learning (RL) framework that progressively optimizes reasoning policies via stage-wise reward shaping. AutoThink learns to invoke explicit reasoning only when necessary, while defaulting to succinct responses for simpler tasks. Experiments on five mainstream mathematical benchmarks demonstrate that AutoThink achieves favorable accuracy–efficiency trade-offs compared to recent prompting and RL-based pruning methods. It can be seamlessly integrated into any R1-style model, including both distilled and further fine-tuned variants. Notably, AutoThink improves relative accuracy by 6.4% while reducing token usage by 52% on DeepSeek-R1-Distill-Qwen-1.5B, establishing a scalable and adaptive reasoning paradigm for LRMs.
Project Page: https://github.com/ScienceOne-AI/AutoThink.
@@ 1 Introduction
Recently, reasoning-focused Large Language Models (LLMs), also referred to as Large Reasoning Models (LRMs) [37], have demonstrated remarkable progress in solving complex reasoning tasks. Particularly, DeepSeek-R1 [7] uses only outcome-based feedback and incentivizes explicit reasoning capabilities through reinforcement learning (RL) with verifiable rewards. DeepSeek-R1 and its distilled models typically follow the <think> and <answer> format, where the <think> process generates explicit, step-by-step reasoning sequences to support obtaining a final answer during the <answer> phase. We refer to models that follow this Chain of Thought (CoT) [33] prompting scheme as R1-style models. The explicit thinking process, which enables self-reflection, backtracking, and validation, is widely regarded as essential for enhancing reasoning accuracy. Arising from this understanding, a popular paradigm has emerged that improves solution quality by increasing thinking token allocation during inference-time reasoning [15, 36]. However, this paradigm introduces a major bottleneck: excessive thinking token generation leads to high computational cost and latency, raising the overthinking phenomenon, where many reasoning steps are redundant or inefficient [28, 12].
To mitigate overthinking, recent efforts have explored hybrid reasoning and concise reasoning strategies. In the industry, Claude 3.7 Sonnet [2] introduces a controllable reasoning framework that allows the model to switch between standard and extended reasoning modes. Similarly, Qwen3 [29] proposes a thinking control scheme with a "thinking" mode (slow thinking) and a "non-thinking" mode (fast thinking), and provides users with the flexibility to choose whether the model should engage in reasoning or not. In the academic community, parallel research has focused on designing prompt-guided efficient reasoning [38, 23] or training pruning-based models to achieve concise reasoning [10, 6, 39]. While promising, these approaches either rely on manually predefined modes or uniformly prune reasoning steps, which may degrade performance on harder instances. A fundamental question then arises to address the overthinking issue:
@@ 2.1 A Surprising Effect of Minimal Prompt Modification
Recent efforts on concise reasoning aim to eliminate unnecessary thought, either via prompting that explicitly bypasses thinking [18], or RL-based training that penalizes long outputs [6]. While effective at shortening responses, these methods enforce uniform brevity regardless of problem complexity. Rather than compressing by default, we pose a subtler question:
@@ 3.1 Stage 1: Preventing Mode Collapse by Batch Reward Balance
While this reward structure encourages difficulty-aware behavior, it causes instability during early training. The model may collapse into a degenerate policy, either always thinking or always skipping, depending on which yields a higher expected reward in the short term. This limits exploration and hinders later optimization. To mitigate this, we introduce batch-level reward balancing:
@@ 3.2 Stage 2: Reinforcing Reliable Behavior within Dual Modes
To allow the model to refine its behavior without external constraints, we remove the batch-level balancing used in the previous stage and allow free evolution of the reasoning policy. The reward is set directly to the naive definition:
@@ 4.1 Setup
AutoThink-Stage2 92.2 38.5 56.2 57.1 83.7 65.5 2519 2980 5797 8676 4925 4979 7.2
AutoThink-Stage3 91.2 38.2 56.4 54.8 83.3 64.8 2146 2838 5498 8051 4645 4635 3.2
@@ 4.4 In-Depth Behavioral and Efficiency Analysis
We further conduct additional analyses on more base models, reasoning behaviors, training cost, and case studies. Details are presented in Appendix B due to space limitations.
@@ Appendix A Additional Prompt Variants
Appendix B Extended Experimental Results
B.1 Additional Results on Skywork-OR1-Math-7B
To further assess the generality of our method, we apply the AutoThink framework to Skywork-OR1-Math-7B [^1] https://huggingface.co/Skywork/Skywork-OR1-Math-7B, a state-of-the-art 7B model that achieves strong performance on mathematical reasoning tasks. Pretrained and fine-tuned with rule-based reinforcement learning on math and code tasks, this model represents one of the strongest 7B-scale math solvers. As shown in Table 5, the Ellipsis Prompt has limited effect on this highly optimized model, inducing only a marginal proportion of no-thinking responses, indicating reduced prompt sensitivity due to its deterministic reasoning policy.
Despite the limited prompt sensitivity, Stage 1 training with batch-level contrastive signals effectively captures and amplifies the model’s latent no-thinking behavior, enabling more balanced reasoning patterns to emerge. Subsequent Stages 2 and 3 progressively refine this behavior. The full AutoThink framework is applied sequentially over three stages, trained for 600, 500, and 30 steps, respectively. Notably, the final stage achieves a nearly 60% reduction in reasoning tokens (from 9053 to 3974), while preserving task accuracy with less than a 2% degradation compared to the standard prompting baseline. These lightweight training phases are sufficient to induce substantial improvements in efficiency, even on strong pretrained models like Skywork-OR1-Math-7B.
Table 5: Accuracy and Token Usage Comparison on Skywork-OR1-Math-7B.
Method Accuracy (%) Token Usage
MATH Minerva Olympiad AIME24 AMC23 AVG MATH Minerva Olympiad AIME24 AMC23 AVG
Base Model: Skywork-OR1-Math-7B
Standard Prompt 94.0 41.2 62.1 67.1 88.3 70.5 4669 7402 10102 14242 8849 9053
No-Thinking Prompt 85.5 26.1 48.1 45.6 68.2 54.7 1033 775 2982 6416 2402 2722
Ellipsis Prompt 94.0 41.8 61.8 69.1 88.0 70.9 4542 7399 10093 13813 8819 8933
AutoThink-Stage1 92.9 36.9 59.3 65.4 86.6 68.2 1894 2177 4616 7068 4074 3966
AutoThink-Stage2 94.0 40.0 62.3 63.1 88.9 69.7 2298 3247 5437 8091 4521 4719
AutoThink-Stage3 93.1 38.8 61.1 62.7 88.2 68.8 1768 2287 4622 7372 3820 3974
We further visualize the training dynamics of Stage 1 in Figure 9, including the proportion of thinking responses, as well as the response length and accuracy stratified by thinking versus no-thinking behaviors. At early stages, almost all responses involve explicit reasoning. However, batch-wise balancing gradually promotes the emergence of no-thinking behavior. A clear modality shift occurs between steps 100 and 200, marked by a sharp increase in no-thinking responses, which directly contributes to the reduction in average response length. To explicitly encourage a balanced distribution between thinking and no-thinking responses throughout training, we set the target balance ratio $\gamma=0.5$ in each of the three stages. Interestingly, while the accuracy of thinking responses slightly decreases during this phase, the overall accuracy continues to improve. This divergence suggests that the model is learning to skip unnecessary reasoning on simpler problems, thereby increasing both efficiency and decision quality through adaptive control over its reasoning mode.
Figure 9: Training Curves of Skywork-OR1-Math-7B on Stage1.
B.2 Reasoning Behavior Profiling
@@
B.3 Training Cost Comparison
@@
B.4 Case Study
Figure 11,12,13 presents some examples comparing four prompting strategies. For easy problems, AutoThink produces correct answers without explicit reasoning, reflecting effective fast thinking. For medium problems, it may activate both reasoning modes, with thinking and no-thinking responses potentially coexisting. For hard problems, the model engages in deeper, slower reasoning, demonstrating iterative understanding and self-verification before arriving at the correct solution. These observations demonstrate how AutoThink adapts its reasoning to problem difficulty, balancing efficiency and reliability through dynamic control of reasoning depth.